# Tracial states on groupoid $C^{*}$-algebras and essential freeness

### v2

@@ 2.3. Tracial states
Given $\varepsilon>0$, take open $U\subseteq\mathcal{G}^{(0)}$ containing the closure of $\mathrm{s}(\mathrm{supp}f)\cap\mathrm{Fix}(\alpha_{B})$ and open $V\supseteq\overline{U}$ such that $\mu(V)<\varepsilon$. Then we take a $\rho_{\varepsilon}\in C_{c}(\mathcal{G}^{(0)})$ such that $0\leq\rho_{\varepsilon}\leq 1$, $\rho_{\varepsilon}|_{\mathrm{s}(\mathrm{supp}f)\setminus V}\equiv 1$ and $\rho_{\varepsilon}|_{\overline{U}}\equiv 0$. Let us consider $f\ast\rho_{\varepsilon}$, which has support in $B$ by Lemma 2.2. Then
@@
This implies that $\mathrm{supp}(f\ast\rho_{\varepsilon})\cap\{\gamma\in B:\mathrm{s}(\gamma)=\mathrm{r}(\gamma)\}=\emptyset$. Otherwise, there exists $\gamma\in\mathrm{supp}(f\ast\rho_{\varepsilon})\cap\{\gamma\in B:\mathrm{s}(\gamma)=\mathrm{r}(\gamma)\}$. If we take a net $\{\gamma_{\lambda}\}_{\lambda}$ in $(\mathrm{s}|_{B})^{-1}(U)$ which converges to $\gamma$ such that $(f\ast\rho_{\varepsilon})(\gamma_{\lambda})\neq 0$, then Lemma 2.2 implies that $\rho_{\varepsilon}(\mathrm{s}(\gamma_{\lambda}))\neq 0$ for each $\lambda$. We reach a contradiction since $\mathrm{s}(\gamma_{\lambda})\in U$ and $\rho_{\varepsilon}|_{U}\equiv 0$. Therefore, the same analysis in the second paragraph above shows that $\tau(f\ast\rho_{\varepsilon})=0$.
Now take $\eta\in C_{c}(B_{0})$ such that $\eta|_{\overline{B}}\equiv 1$ and $||\eta||_{\infty}=1$, and define $f_{0}\in C_{c}(\mathcal{G}^{(0)})$ by $f_{0}(x)=f\circ(\mathrm{s}|_{B_{0}})^{-1}(x)$ if $x\in\mathrm{s}(B_{0})$ and zero otherwise. As $\mathrm{supp}f\subseteq B_{0}$, $f_{0}$ is a continuous function on $\mathcal{G}^{(0)}$ such that $\mathrm{supp}f_{0}=\mathrm{s}(\mathrm{supp}f)$. Then Lemma 2.2 implies that
@@
$|\tau(f)|$ $=|\tau(f)-\tau(f\ast\rho_{\varepsilon})|$
$=|\tau(\eta\ast(f_{0}-f_{0}\cdot\rho_{\varepsilon}))|$
$\leq\tau(|f_{0}-f_{0}\cdot\rho_{\varepsilon}|)^{1/2}\cdot\tau(\eta\ast|f_{0}-f_{0}\cdot\rho_{\varepsilon}|\ast\eta^{*})^{1/2}$
$\leq\tau(|f_{0}-f_{0}\cdot\rho_{\varepsilon}|)\cdot||\eta^{*}\ast\eta||_{\infty}$
$=\tau(|f_{0}-f_{0}\cdot\rho_{\varepsilon}|),$
where we use the Cauchy-Schwarz inequality $|\tau(y^{*}x)|^{2}\leq\tau(x^{*}x)\cdot\tau(y^{*}y)$ for any $x,y\in C^{*}_{\nu}(\mathcal{G})$ in the first inequality, and the tracial property of $\tau$ in the second inequality. Finally, from $\mu(V)<\epsilon$ we have that
$\tau(|f_{0}-f_{0}\cdot\rho_{\varepsilon}|)=\int_{\mathcal{G}^{(0)}}|f_{0}|\cdot|1-\rho_{\varepsilon}|\mathrm{d}\mu=\int_{\mathrm{s}(\mathrm{supp}f)\cap V}|f_{0}|\cdot|1-\rho_{\varepsilon}|\mathrm{d}\mu\leq\varepsilon\cdot\|f_{0}\|_{\infty},$
@@
If we assume that $\mathcal{G}^{(0)}$ is compact then the convex set $M(\mathcal{G})$ of invariant probability Radon measures on $\mathcal{G}^{(0)}$ and the tracial state space $T(C^{*}_{\nu}(\mathcal{G}))$ of $C^{*}_{\nu}(\mathcal{G})$ are both compact in the ${}^{\ast}$-topology. Then the following corollary of Proposition 3.4 has generalised [1, Proposition 2.2], because almost finite ample groupoids with compact unit space are always essentially free by [23, Remark 6.6]:
@@
In the following, we would like to study Question 1.1. For that we consider an auxiliary trace $\tau^{\mathrm{Fix}}_{\mu}:C_{c}(\mathcal{G})\to\mathbb{C}$ associated to a given invariant probability Radon measure $\mu$ on $\mathcal{G}^{(0)}$. The key point is that $\tau^{\mathrm{Fix}}_{\mu}$ reveals the complete information of essential freeness with respect to $\mu$ (see Lemma 3.14 for details). Since its construction is a bit complicated, we divide it into several steps.
@@
Now we are ready to give a general definition of $\tau^{\mathrm{Fix}}_{\mu}$ on $C_{c}(\mathcal{G})$.
Definition 3.8.
Let $\mathcal{G}$ be a locally compact Hausdorff and étale groupoid, and $\mu$ be an invariant probability Radon measure on $\mathcal{G}^{(0)}$. The associated fixed point trace is the linear map $\tau^{\mathrm{Fix}}_{\mu}:C_{c}(\mathcal{G})\to\mathbb{C}$ defined as follows: For $g\in C_{c}(\mathcal{G})$ with $g=\sum_{i=1}^{n}g_{i}$ such that each $g_{i}\in C_{c}(B_{i})$ for some pre-compact open bisection $B_{i}$. Then we define
where $\tau^{\mathrm{Fix}}_{\mu}(g_{i})$ is given in (3.1).
In the next lemma, we show that $\tau^{\mathrm{Fix}}_{\mu}$ in the above definition is well-defined:
@@
Since $K:=\bigcup_{i=1}^{n}\mathrm{supp}(g_{i})\cup\bigcup_{j=1}^{m}\mathrm{supp}(h_{j})$ is compact, we can take a finite open cover $\mathcal{U}=\{U_{k}:k=1,\cdots,N\}$ of $K$ such that each $U_{k}$ is a pre-compact open bisection. Then we take a partition of unity $\{\rho_{k}:k=1,\cdots,N\}$ subordinate to $\mathcal{U}$ such that $\sum_{k=1}^{N}\rho_{k}\equiv 1$ on $K$. In particular, we have $g_{i}=\sum_{k=1}^{N}(\rho_{k}g_{i})$ for any $i=1,\cdots,n$. As both $\mathrm{supp}(g_{i})$ and $\mathrm{supp}(\rho_{k}g_{i})$ are contained in $B_{i}$, it follows from Lemma 3.7 that
(3.3) $\tau^{\mathrm{Fix}}_{\mu}(g_{i})=\int_{\mathrm{Fix}(\alpha_{B_{i}})}(g_{i})_{B_{i}}\mathrm{d}\mu\quad\text{and}\quad\tau^{\mathrm{Fix}}_{\mu}(\rho_{k}g_{i})=\int_{\mathrm{Fix}(\alpha_{B_{i}})}(\rho_{k}g_{i})_{B_{i}}\mathrm{d}\mu.$
For each $x\in\mathrm{s}(B_{i})$, we have $(\rho_{k}g_{i})_{B_{i}}(x)=\rho_{k}((s|_{B_{i}})^{-1}(x))\cdot g_{i}((s|_{B_{i}})^{-1}(x))$. Hence, we obtain $\sum_{k=1}^{N}(\rho_{k}g_{i})_{B_{i}}=(g_{i})_{B_{i}}$, which together with (3.3) implies that
$\tau^{\mathrm{Fix}}_{\mu}(g_{i})=\int_{\mathrm{Fix}(\alpha_{B_{i}})}(g_{i})_{B_{i}}\mathrm{d}\mu=\sum_{k=1}^{N}\int_{\mathrm{Fix}(\alpha_{B_{i}})}(\rho_{k}g_{i})_{B_{i}}\mathrm{d}\mu=\sum_{k=1}^{N}\tau^{\mathrm{Fix}}_{\mu}(\rho_{k}g_{i}).$
@@
(3.4) $\sum_{i=1}^{n}\tau^{\mathrm{Fix}}_{\mu}(g_{i})=\sum_{i=1}^{n}\sum_{k=1}^{N}\tau^{\mathrm{Fix}}_{\mu}(\rho_{k}g_{i})=\sum_{k=1}^{N}\sum_{i=1}^{n}\tau^{\mathrm{Fix}}_{\mu}(\rho_{k}g_{i}).$
@@
(3.5) $\sum_{j=1}^{m}\tau^{\mathrm{Fix}}_{\mu}(h_{j})=\sum_{k=1}^{N}\sum_{j=1}^{m}\tau^{\mathrm{Fix}}_{\mu}(\rho_{k}h_{j}).$
For a fixed $k\in\{1,\cdots,N\}$, using a similar argument as above we have
(3.6) $\sum_{i=1}^{n}\tau^{\mathrm{Fix}}_{\mu}(\rho_{k}g_{i})=\tau^{\mathrm{Fix}}_{\mu}\big{(}\sum_{i=1}^{n}\rho_{k}g_{i}\big{)}=\tau^{\mathrm{Fix}}_{\mu}(\rho_{k}g)=\tau^{\mathrm{Fix}}_{\mu}\big{(}\sum_{j=1}^{m}\rho_{k}h_{j}\big{)}=\sum_{j=1}^{m}\tau^{\mathrm{Fix}}_{\mu}(\rho_{k}h_{j}).$
@@
Assume that $a=\sum_{i=1}^{n}g_{i}$ such that each $g_{i}\in C_{c}(B_{i})$ for some pre-compact open bisection $B_{i}$. Then we have $a^{*}a=\sum_{i,j=1}^{n}g_{i}^{*}\ast g_{j}$ and $aa^{*}=\sum_{i,j=1}^{n}g_{j}\ast g_{i}^{*}$. Note that for any $\gamma\in\mathcal{G}$ and $i,j=1,\cdots,n$, we have
@@
It is clear that $B_{j}\cdot B_{i}^{-1}$ is precompact, and it follows from Lemma 2.1 that $B_{j}\cdot B_{i}^{-1}$ is also an open bisection. Moreover, for any $\gamma\in B_{j}\cdot B_{i}^{-1}$ there exist unique $\beta\in B_{j}$ and $\alpha\in B_{i}^{-1}$ such that $\gamma=\beta\alpha$, and hence $(g_{j}\ast g_{i}^{*})(\gamma)=g_{j}(\beta)\cdot\overline{g_{i}(\alpha^{-1})}$.
@@
Thus, it is easy to check that $\alpha_{B_{j}\cdot B_{i}^{-1}}=\alpha_{B_{j}}\circ\alpha_{B_{i}}^{-1}$. Moreover, if $x\in\mathrm{Fix}(\alpha_{B_{j}}\circ\alpha_{B_{i}}^{-1})$ then we take $\gamma\in B_{j}\cdot B_{i}^{-1}$ such that $\mathrm{s}(\gamma)=x$. Writing $\gamma=\beta\alpha^{-1}$ as before for $\beta\in B_{j}$ and $\alpha\in B_{i}$, then $\mathrm{s}(\alpha)=\mathrm{s}(\beta)$ and $\mathrm{r}(\alpha)=x=\alpha_{B_{j}}\circ\alpha_{B_{i}}^{-1}(x)=\alpha_{B_{j}}(\mathrm{s}(\alpha))=\alpha_{B_{j}}(\mathrm{s}(\beta))=\mathrm{r}(\beta)=\mathrm{r}(\gamma)$. So for $x\in\mathrm{Fix}(\alpha_{B_{j}}\circ\alpha_{B_{i}}^{-1})$ we obtain
@@
$\tau^{\mathrm{Fix}}_{\mu}(aa^{*})=\sum_{i,j=1}^{n}\int_{\mathrm{Fix}(\alpha_{B_{j}}\circ\alpha_{B_{i}}^{-1})}(g_{j})^{B_{j}}(x)\cdot\overline{(g_{i})^{B_{i}}(x)}\mathrm{d}\mu(x)=\int_{\mathcal{G}^{(0)}}\sum_{i,j=1}^{n}\chi_{B_{ij}}(g_{j})^{B_{j}}(x)\cdot\overline{(g_{i})^{B_{i}}(x)}\mathrm{d}\mu(x),$
@@
$F(x):=\sum_{i,j=1}^{n}\chi_{B_{ij}}(g_{j})^{B_{j}}(x)\cdot\overline{(g_{i})^{B_{i}}(x)}\quad\text{for}\quad x\in\mathcal{G}^{(0)}.$
It suffices to show that $F(x)\geq 0$ for each $x\in\mathcal{G}^{(0)}$. Fix an $x\in\mathcal{G}^{(0)}$, and let $I:=\{i\in\{1,\cdots,n\}:x\in\mathrm{r}(B_{i})\}$. Decompose $I=I_{1}\sqcup I_{2}\sqcup\cdots\sqcup I_{N}$ such that for any $i,j\in I_{k}$ (where $k=1,\cdots,N$) we have $\alpha_{B_{i}}^{-1}(x)=\alpha_{B_{j}}^{-1}(x)$, and for any $i\in I_{k}$ and $j\in I_{l}$ where $k\neq l$ and $k,l=1,\cdots,N$ we have $\alpha_{B_{i}}^{-1}(x)\neq\alpha_{B_{j}}^{-1}(x)$. Hence, we have
$F(x)=\sum_{k=1}^{N}\sum_{i,j\in I_{k}}(g_{j})^{B_{j}}(x)\cdot\overline{(g_{i})^{B_{i}}(x)}=\sum_{k=1}^{N}\big{|}\sum_{i\in I_{k}}(g_{i})^{B_{i}}(x)\big{|}^{2}\geq 0.$
So we have completed the proof. ∎
@@
By Lemma 3.10, $\tau_{\mu}^{\mathrm{Fix}}|_{C_{c}(\mathcal{G}^{(0)})}$ can be extended to a tracial state on $C_{0}(\mathcal{G}^{(0)})$ (with the same notation). Hence, we have $\lim_{i\in I}\tau^{\mathrm{Fix}}_{\mu}(u_{i})=\|\tau_{\mu}^{\mathrm{Fix}}|_{C_{0}(\mathcal{G}^{(0)})}\|=1$. ∎
@@
“(1) $\Rightarrow$ (2)”: The first half sentence follows from Proposition 3.4. For the second statement, Lemma 3.14 shows that $\tau^{\mathrm{Fix}}_{\mu}(a)=\tau_{\mu}(a)$ for $a\in C_{c}(\mathcal{G})$. Since $\tau_{\mu}$ can be extended to a tracial state on $C^{*}_{\nu}(\mathcal{G})$, we conclude (2).
@@
Recall that a groupoid $\mathcal{G}$ has the weak containment property if $C^{*}(\mathcal{G})\cong C^{*}_{r}(\mathcal{G})$ canonically [^3] Amenable groupoids have the weak containment property, but there are also non-amenable groupoids with the weak containment property [34].. In this case, we obtain the following:

### v3

@@ 2.3. Tracial states
Since $\mu$ is a Radon measure, it is outer regular. Hence given $\varepsilon>0$, we can take an open set $U\subseteq\mathcal{G}^{(0)}$ containing the closure of $\mathrm{s}(\mathrm{supp}f)\cap\mathrm{Fix}(\alpha_{B})$ and an open set $V\supseteq\overline{U}$ such that $\mu(V)<\varepsilon$. Since $\mathrm{s}(\mathrm{supp}f)$ is compact and $V$ is open, $\mathrm{s}(\mathrm{supp}f)\setminus V$ is also compact. Hence we can take $\rho_{\varepsilon}\in C_{c}(\mathcal{G}^{(0)})$ such that $0\leq\rho_{\varepsilon}\leq 1$, $\rho_{\varepsilon}|_{\mathrm{s}(\mathrm{supp}f)\setminus V}\equiv 1$ and $\rho_{\varepsilon}|_{\overline{U}}\equiv 0$.
We aim to apply the argument of the second paragraph above to the function $f\ast\rho_{\varepsilon}$, which has support in $B$ by Lemma 2.2. Hence we must first show that $\mathrm{supp}(f\ast\rho_{\varepsilon})\cap\{\gamma\in B:\mathrm{s}(\gamma)=\mathrm{r}(\gamma)\}=\emptyset$. Note that
@@
Assume that there exists $\gamma\in\mathrm{supp}(f\ast\rho_{\varepsilon})\cap\{\gamma\in B:\mathrm{s}(\gamma)=\mathrm{r}(\gamma)\}$. Since $\gamma\in(\mathrm{s}|_{B})^{-1}(U)\cap\mathrm{supp}(f\ast\rho_{\varepsilon})$, we can choose a net $\{\gamma_{\lambda}\}_{\lambda}$ in $(\mathrm{s}|_{B})^{-1}(U)$ converging to $\gamma$ such that $(f\ast\rho_{\varepsilon})(\gamma_{\lambda})\neq 0$, then Lemma 2.2 implies that $\rho_{\varepsilon}(\mathrm{s}(\gamma_{\lambda}))\neq 0$ for each $\lambda$. We reach a contradiction since $\mathrm{s}(\gamma_{\lambda})\in U$ and $\rho_{\varepsilon}|_{U}\equiv 0$. Therefore, the same analysis in the second paragraph of this proof shows that $\tau(f\ast\rho_{\varepsilon})=0$.
Take $\eta\in C_{c}(B_{0})$ such that $\eta|_{\overline{B}}\equiv 1$ and $||\eta||_{\infty}=1$, and define $f_{0}\in C_{c}(\mathcal{G}^{(0)})$ by $f_{0}(x)=f\circ(\mathrm{s}|_{B_{0}})^{-1}(x)$ if $x\in\mathrm{s}(B_{0})$ and zero otherwise. As $\mathrm{supp}f\subseteq B_{0}$, $f_{0}$ is a continuous function on $\mathcal{G}^{(0)}$ such that $\mathrm{supp}f_{0}=\mathrm{s}(\mathrm{supp}f)$. Then Lemma 2.2 implies that
@@
(3.1) $|\tau(f)|=|\tau(f)-\tau(f\ast\rho_{\varepsilon})|=|\tau(\eta\ast(f_{0}-f_{0}\cdot\rho_{\varepsilon}))|.$
Since $f_{0}-f_{0}\cdot\rho_{\varepsilon}$ is point-wise non-negative, we write
$\eta\ast(f_{0}-f_{0}\cdot\rho_{\varepsilon})=\left(\eta\ast(f_{0}-f_{0}\cdot\rho_{\varepsilon})^{1/2}\right)\ast(f_{0}-f_{0}\cdot\rho_{\varepsilon})^{1/2}.$
Using the Cauchy–Schwarz inequality $|\tau(y^{*}x)|^{2}\leq\tau(x^{*}x)\cdot\tau(y^{*}y)$ for $x=(f_{0}-f_{0}\cdot\rho_{\varepsilon})^{1/2}$ and $y^{*}=\eta\ast(f_{0}-f_{0}\cdot\rho_{\varepsilon})^{1/2}$, we obtain
(3.2) $|\tau(\eta\ast(f_{0}-f_{0}\cdot\rho_{\varepsilon}))|^{2}\leq\tau(f_{0}-f_{0}\cdot\rho_{\varepsilon})\cdot\tau(\eta\ast(f_{0}-f_{0}\cdot\rho_{\varepsilon})\ast\eta^{*}).$
Using properties of tracial states, we have
(3.3) $\tau(\eta\ast(f_{0}-f_{0}\cdot\rho_{\varepsilon})\ast\eta^{*})=\tau((f_{0}-f_{0}\cdot\rho_{\varepsilon})^{1/2}\ast\eta^{*}\ast\eta\ast(f_{0}-f_{0}\cdot\rho_{\varepsilon})^{1/2})\leq\tau(f_{0}-f_{0}\cdot\rho_{\varepsilon}),$
where we use $\|\eta\|_{\infty}=1$ for the second inequality. Combining (3.1), (3.2) and (3.3), we obtain
$|\tau(f)|\leq\tau(f_{0}-f_{0}\cdot\rho_{\varepsilon}).$
Finally, from $\mu(V)<\epsilon$ we have that
$\tau(f_{0}-f_{0}\cdot\rho_{\varepsilon})=\int_{\mathcal{G}^{(0)}}f_{0}\cdot(1-\rho_{\varepsilon})\mathrm{d}\mu=\int_{\mathrm{s}(\mathrm{supp}f)\cap V}f_{0}\cdot(1-\rho_{\varepsilon})\mathrm{d}\mu\leq\varepsilon\cdot\|f_{0}\|_{\infty},$
@@
If we assume that $\mathcal{G}^{(0)}$ is compact then the convex set $M(\mathcal{G})$ of invariant probability Radon measures on $\mathcal{G}^{(0)}$ and the tracial state space $T(C^{*}_{\nu}(\mathcal{G}))$ of $C^{*}_{\nu}(\mathcal{G})$ are both compact in the weak-topology. Then the following corollary of Proposition 3.4 has generalised [1, Proposition 3.1], because almost finite ample groupoids with compact unit space are always essentially free by [23, Remark 6.6]:
@@
In the following, we would like to study Question 1.1 for a locally compact Hausdorff and étale groupoid such that $\mathcal{G}^{(0)}$ is not necessarily compact. For that we consider an auxiliary trace $\tau^{\mathrm{Fix}}_{\mu}:C_{c}(\mathcal{G})\to\mathbb{C}$ associated to a given invariant probability Radon measure $\mu$ on $\mathcal{G}^{(0)}$. The key point is that $\tau^{\mathrm{Fix}}_{\mu}$ reveals the complete information of essential freeness with respect to $\mu$ (see Lemma 3.14 for details). Since its construction is a bit complicated, we divide it into several steps.
@@
Moreover, we have the following:
@@
Since $K:=\bigcup_{i=1}^{n}\mathrm{supp}(g_{i})\cup\bigcup_{j=1}^{m}\mathrm{supp}(h_{j})$ is compact, we can take a finite open cover $\mathcal{U}=\{U_{k}:k=1,\dotsc,N\}$ of $K$ such that each $U_{k}$ is a pre-compact open bisection. Then we take a partition of unity $\{\rho_{k}:k=1,\dotsc,N\}$ subordinate to $\mathcal{U}$ such that $\sum_{k=1}^{N}\rho_{k}\equiv 1$ on $K$. In particular, we have $g_{i}=\sum_{k=1}^{N}(\rho_{k}\cdot g_{i})$ for any $i=1,\dotsc,n$, where $\rho_{k}\cdot g_{i}$ means the point-wise product as in Section 2.2. As both $\mathrm{supp}(g_{i})$ and $\mathrm{supp}(\rho_{k}\cdot g_{i})$ are contained in $B_{i}$, it follows from Lemma 3.7 that
(3.5) $\tau^{\mathrm{Fix}}_{\mu}(g_{i})=\int_{\mathrm{Fix}(\alpha_{B_{i}})}(g_{i})_{B_{i}}\mathrm{d}\mu\quad\text{and}\quad\tau^{\mathrm{Fix}}_{\mu}(\rho_{k}\cdot g_{i})=\int_{\mathrm{Fix}(\alpha_{B_{i}})}(\rho_{k}\cdot g_{i})_{B_{i}}\mathrm{d}\mu.$
For each $x\in\mathrm{s}(B_{i})$, we have $(\rho_{k}\cdot g_{i})_{B_{i}}(x)=\rho_{k}((s|_{B_{i}})^{-1}(x))\cdot g_{i}((s|_{B_{i}})^{-1}(x))$. Hence, we obtain $\sum_{k=1}^{N}(\rho_{k}\cdot g_{i})_{B_{i}}=(g_{i})_{B_{i}}$, which together with (3.5) implies that
$\tau^{\mathrm{Fix}}_{\mu}(g_{i})=\int_{\mathrm{Fix}(\alpha_{B_{i}})}(g_{i})_{B_{i}}\mathrm{d}\mu=\sum_{k=1}^{N}\int_{\mathrm{Fix}(\alpha_{B_{i}})}(\rho_{k}\cdot g_{i})_{B_{i}}\mathrm{d}\mu=\sum_{k=1}^{N}\tau^{\mathrm{Fix}}_{\mu}(\rho_{k}\cdot g_{i}).$
@@
(3.6) $\sum_{i=1}^{n}\tau^{\mathrm{Fix}}_{\mu}(g_{i})=\sum_{i=1}^{n}\sum_{k=1}^{N}\tau^{\mathrm{Fix}}_{\mu}(\rho_{k}\cdot g_{i})=\sum_{k=1}^{N}\sum_{i=1}^{n}\tau^{\mathrm{Fix}}_{\mu}(\rho_{k}\cdot g_{i}).$
@@
(3.7) $\sum_{j=1}^{m}\tau^{\mathrm{Fix}}_{\mu}(h_{j})=\sum_{k=1}^{N}\sum_{j=1}^{m}\tau^{\mathrm{Fix}}_{\mu}(\rho_{k}\cdot h_{j}).$
For a fixed $k\in\{1,\dotsc,N\}$, using a similar argument as above we have
(3.8) $\sum_{i=1}^{n}\tau^{\mathrm{Fix}}_{\mu}(\rho_{k}\cdot g_{i})=\tau^{\mathrm{Fix}}_{\mu}\big{(}\sum_{i=1}^{n}\rho_{k}\cdot g_{i}\big{)}=\tau^{\mathrm{Fix}}_{\mu}(\rho_{k}\cdot g)=\tau^{\mathrm{Fix}}_{\mu}\big{(}\sum_{j=1}^{m}\rho_{k}\cdot h_{j}\big{)}=\sum_{j=1}^{m}\tau^{\mathrm{Fix}}_{\mu}(\rho_{k}\cdot h_{j}).$
@@
Lemma 3.8 allows us to give the following definition of $\tau^{\mathrm{Fix}}_{\mu}$ on $C_{c}(\mathcal{G})$.
Definition 3.9.
Let $\mathcal{G}$ be a locally compact Hausdorff and étale groupoid, and $\mu$ be an invariant probability Radon measure on $\mathcal{G}^{(0)}$. The associated fixed point trace is the linear map $\tau^{\mathrm{Fix}}_{\mu}:C_{c}(\mathcal{G})\to\mathbb{C}$ defined as follows: Suppose that $g\in C_{c}(\mathcal{G})$ with $g=\sum_{i=1}^{n}g_{i}$ such that each $g_{i}\in C_{c}(B_{i})$ for some pre-compact open bisection $B_{i}$. Then we define
where $\tau^{\mathrm{Fix}}_{\mu}(g_{i})$ is given in (3.4).
@@
Assume that $a=\sum_{i=1}^{n}g_{i}$ such that each $g_{i}\in C_{c}(B_{i})$ for some pre-compact open bisection $B_{i}$. Then we have $a^{*}a=\sum_{i,j=1}^{n}g_{i}^{*}\ast g_{j}$ and $aa^{*}=\sum_{i,j=1}^{n}g_{j}\ast g_{i}^{*}$. Note that for any $\gamma\in\mathcal{G}$ and $i,j=1,\dotsc,n$, we have
@@
Since multiplication and inversion are continuous and $\mathcal{G}$ is Hausdorff, the set $B_{j}\cdot B_{i}^{-1}$ is precompact, and it follows from Lemma 2.1 that $B_{j}\cdot B_{i}^{-1}$ is also an open bisection. Moreover, for any $\gamma\in B_{j}\cdot B_{i}^{-1}$ there exist unique $\beta\in B_{j}$ and $\alpha\in B_{i}^{-1}$ such that $\gamma=\beta\alpha$, and hence $(g_{j}\ast g_{i}^{*})(\gamma)=g_{j}(\beta)\cdot\overline{g_{i}(\alpha^{-1})}$.
@@
Thus by definition, a direct calculation shows that $\alpha_{B_{j}\cdot B_{i}^{-1}}=\alpha_{B_{j}}\circ\alpha_{B_{i}}^{-1}$. Moreover, if $x\in\mathrm{Fix}(\alpha_{B_{j}}\circ\alpha_{B_{i}}^{-1})$ then we take $\gamma\in B_{j}\cdot B_{i}^{-1}$ such that $\mathrm{s}(\gamma)=x$. Writing $\gamma=\beta\alpha^{-1}$ for $\beta\in B_{j}$ and $\alpha\in B_{i}$, then $\mathrm{s}(\alpha)=\mathrm{s}(\beta)$ and $\mathrm{r}(\alpha)=x=\alpha_{B_{j}}\circ\alpha_{B_{i}}^{-1}(x)=\alpha_{B_{j}}(\mathrm{s}(\alpha))=\alpha_{B_{j}}(\mathrm{s}(\beta))=\mathrm{r}(\beta)=\mathrm{r}(\gamma)$. So for $x\in\mathrm{Fix}(\alpha_{B_{j}}\circ\alpha_{B_{i}}^{-1})$ we obtain
@@
$\tau^{\mathrm{Fix}}_{\mu}(aa^{*})$ $=\sum_{i,j=1}^{n}\int_{\mathrm{Fix}(\alpha_{B_{j}}\circ\alpha_{B_{i}}^{-1})}(g_{j})^{B_{j}}(x)\cdot\overline{(g_{i})^{B_{i}}(x)}\mathrm{d}\mu(x)$
$=\int_{\mathcal{G}^{(0)}}\sum_{i,j=1}^{n}\chi_{B_{ij}}(x)\cdot(g_{j})^{B_{j}}(x)\cdot\overline{(g_{i})^{B_{i}}(x)}\mathrm{d}\mu(x),$
@@
$F(x):=\sum_{i,j=1}^{n}\chi_{B_{ij}}(x)\cdot(g_{j})^{B_{j}}(x)\cdot\overline{(g_{i})^{B_{i}}(x)}\quad\text{for}\quad x\in\mathcal{G}^{(0)}.$
It suffices to show that $F(x)\geq 0$ for each $x\in\mathcal{G}^{(0)}$. Fix an $x\in\mathcal{G}^{(0)}$, and let $I:=\{i\in\{1,\dotsc,n\}:x\in\mathrm{r}(B_{i})\}$. Note that there is an equivalence relation on elements of $I$ given by $i\sim j$ if and only if $x\in B_{ij}$. Hence we can decompose $I=I_{1}\sqcup I_{2}\sqcup\dotsb\sqcup I_{N}$ such that for any $i,j\in I_{k}$ (where $k=1,\dotsc,N$) we have $\alpha_{B_{i}}^{-1}(x)=\alpha_{B_{j}}^{-1}(x)$, and for any $i\in I_{k}$ and $j\in I_{l}$ where $k\neq l$ and $k,l=1,\dotsc,N$ we have $\alpha_{B_{i}}^{-1}(x)\neq\alpha_{B_{j}}^{-1}(x)$. Hence, we have
$F(x)=\sum_{k=1}^{N}\sum_{i,j\in I_{k}}(g_{j})^{B_{j}}(x)\cdot\overline{(g_{i})^{B_{i}}(x)}=\sum_{k=1}^{N}\big{|}\sum_{i\in I_{k}}(g_{i})^{B_{i}}(x)\big{|}^{2}\geq 0,$
where we use the elementary calculation $|\sum_{i=1}^{n}\lambda_{i}|^{2}=\sum_{i,j=1}^{n}\lambda_{i}\overline{\lambda_{j}}$ for $\lambda_{1},\ldots,\lambda_{n}\in\mathbb{C}$ for the second equality. So we have completed the proof. ∎
@@
By Lemma 3.10, $\tau_{\mu}^{\mathrm{Fix}}|_{C_{c}(\mathcal{G}^{(0)})}$ can be extended to a tracial state on $C_{0}(\mathcal{G}^{(0)})$ (with the same notation). Hence, we have $\lim_{i\in I}\tau^{\mathrm{Fix}}_{\mu}(u_{i})=\|\tau_{\mu}^{\mathrm{Fix}}|_{C_{0}(\mathcal{G}^{(0)})}\|=1$ (see, e.g., [24, Theorem 3.3.3]). ∎
@@
We claim that $\{[\rho_{K}^{1/2}]\}_{K}$ is a Cauchy net in $\mathcal{H}$. In fact, we have
$\left\|[\rho_{K}^{1/2}]-[\rho_{L}^{1/2}]\right\|^{2}_{\mathcal{H}}$ $=\tau^{\mathrm{Fix}}_{\mu}\left((\rho_{K}^{1/2}-\rho_{L}^{1/2})(\rho_{K}^{1/2}-\rho_{L}^{1/2})\right)$
$=\tau^{\mathrm{Fix}}_{\mu}(\rho_{K})+\tau^{\mathrm{Fix}}_{\mu}(\rho_{L})-2\tau^{\mathrm{Fix}}_{\mu}(\rho_{K}^{1/2}\rho_{L}^{1/2}).$
Taking limits for $K$ and $L$, Lemma 3.12 implies that $\tau^{\mathrm{Fix}}_{\mu}(\rho_{K})\to 1$ and $\tau^{\mathrm{Fix}}_{\mu}(\rho_{L})\to 1$. Moreover, Lemma 3.10 shows that
$\tau^{\mathrm{Fix}}_{\mu}(\rho_{K}^{1/2}\rho_{L}^{1/2})=\int_{\mathcal{G}^{(0)}}\rho_{K}^{1/2}\rho_{L}^{1/2}\mathrm{d}\mu\to 1$
since $\mu$ is a probability Radon measure. This concludes the proof that $\{[\rho_{K}^{1/2}]\}_{K}$ is a Cauchy net in $\mathcal{H}$, and hence converges to a unit vector $\xi\in\mathcal{H}$. For $a\in C^{*}(\mathcal{G})$, we define $\tau^{\mathrm{Fix}}_{\mu}(a):=\langle\pi(a)\xi,\xi\rangle$. It is routine to check that $\tau^{\mathrm{Fix}}_{\mu}$ is a tracial state on $C^{*}(\mathcal{G})$ which extends $\tau^{\mathrm{Fix}}_{\mu}:C_{c}(\mathcal{G})\to\mathbb{C}$. ∎
@@
“(1) $\Rightarrow$ (2)”: The first half of the sentence follows from Proposition 3.4. For the second statement, Lemma 3.14 shows that $\tau^{\mathrm{Fix}}_{\mu}(a)=\tau_{\mu}(a)$ for $a\in C_{c}(\mathcal{G})$. Since $\tau_{\mu}$ can be extended to a tracial state on $C^{*}_{\nu}(\mathcal{G})$, we conclude (2).
@@
Recall that a groupoid $\mathcal{G}$ has the weak containment property if $C^{*}(\mathcal{G})\cong C^{*}_{r}(\mathcal{G})$ canonically [^3] Amenable groupoids have the weak containment property, but there are also non-amenable groupoids with the weak containment property [35].. In this case, we obtain the following directly from Corollary 3.13 and Theorem 3.16: